OpenStudy (anonymous):

FAN AND MEDAL. What is the 9th term of the geometric sequence where a1 = -8 and a6 = -8,192?

OpenStudy (anonymous):

A. -327,680 B. -393,216 C. -458,752 D. -524,288

OpenStudy (anonymous):

\[a_n = a_1 q^{n-1}\], so... \[a_6 = a_1 q^{5}\]\[-8.192 = -8 q^{5}\]\[q = 1.00475\] Now, just find the a_9

OpenStudy (michele_laino):

please note that: \[a _{6}=a _{1}*q ^{5}\]whre q is the constant f yr geometric sequence

OpenStudy (anonymous):

Do you understand?

OpenStudy (michele_laino):

oops...of your geometric...

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

The geometric sequency have the form: \[a_n = a_1 q^{n-1}\]

OpenStudy (anonymous):

if we know the ratio "q" and the first element of the sequency "a_1", we can calculate any element of this sequency

OpenStudy (michele_laino):

please substitute your umerical data ino the subsequent relation: \[a _{6}=a _{1}*q ^{5}\] and find q^5, please!

OpenStudy (anonymous):

you know the first element "a_1" and element "a_6", using the equation above, you need to find the ratio of the sequency, for just then, find the value of "a_9" by: \[a_9 = a_1 q^{8}\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I'm actually confused.

OpenStudy (anonymous):

have u ever seen the equation i showed you?

OpenStudy (anonymous):

In what exactly you got confused?

OpenStudy (anonymous):

No, I don't think I've ever seen the equation and I'm just not awesome at math.

OpenStudy (anonymous):

Ok, let's deduce the formula, suppose we have the element one of the sequency \(a_1\) and we what to find the second element of the sequency, then: \(a_2 = a_1*q\), the third element will be \(a_3 = a_2 * q\), and so on... For the n-th element: \(a_n = a_{n-1} * q = a_{n-2}*q^2 = ... = a_1*q^{n-1}\)

OpenStudy (anonymous):

It's multiple choice on a timed test so honestly I'm just gonna guess at this point. Thanks for your help though.

OpenStudy (shaik0124):

q=1.00475 a1 = -8 a9=a1 q^8 a9=-8*(1.00475)^8 slve u will get answer

OpenStudy (anonymous):

Thanks!

OpenStudy (shaik0124):

did u get the answer

OpenStudy (anonymous):

Yeah, I did. It's -524,288 right?

OpenStudy (shaik0124):

whats the value of this (1.00475)^8 ???

OpenStudy (anonymous):

I've already submitted the test so if it isn't right, then oh well.

OpenStudy (shaik0124):

just say whats the value of this (1.00475)^8 ???

OpenStudy (anonymous):

1.0386377874

OpenStudy (anonymous):

I got the right answer so it's all good.

OpenStudy (shaik0124):

after multiplying1.0386377874 with -8 the answer is ??

OpenStudy (anonymous):

There's no point to this anymore, the final answer is -524,288. That's what I needed.

OpenStudy (shaik0124):

ok

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (shaik0124):

i didnt get the answer